Jet 1 can fly from San Francisco to Dallas in 3 hours. Jet 2 can make the same trip in 3.5 hours. If the two jets leave San Francisco at noon, after how many hours is Jet 2 twice as far from Dallas as jet 1?
Answer by scott8148(5891)

the question is asking how long until the FRACTION of the trip remaining for jet 2 is twice the FRACTION remaining for jet 1

let x=hours __ (3.5-x)/3.5=2((3-x)/3)

multiplying by 3(3.5) __ 10.5-3x=21-7x __ 4x=10.5